kjo28050:

A piece of wire 6 cm long weighs 30 grams. What is the weight of an 8 m length of the same wire?A)400 gramsB)425 gramsC)640 gramsD)4000 gram

Vocaloid:

Set up a proportion Mass of wire/length of wire = mass of wire/length of wire Since we have one unit in cm and one unit in m, let’s convert both units to cm 8m = 8 * 100 cm/m = 800cm Plugging in: 30g/6 cm = x g / 800 cm Solve for x